OpenStudy (help!!!!):

36. During a 5-day festival, the number of visitors tripled each day. If the festival opened on a Thursday with 345 visitors, what was the attendance on that Sunday? A. 345 B. 1,035 C. 1,725 D. 3,105 E. 9,315

OpenStudy (anonymous):

Thursday to Friday: 1035. Friday to Saturday: 3105. Saturday to Sunday: *9315*.

OpenStudy (anonymous):

Day 1 (Thursday) = 345 Day 2 (Friday) = 345(3) = 1035 Day 3 (Saturday) = Day2(3) = 3105 Day 4 (Sunday) = Day3(3) = 9315 ...I would say 9315.

OpenStudy (help!!!!):

how did you get the answer, is there a formula for this?

OpenStudy (anonymous):

I just kept multiplying by 3.

OpenStudy (unklerhaukus):

\[\large A_{\text{sunday}}=A_{\text{thursday}}\times3^{|\text{sunday}-\text{thursday}|}\]\[\quad\qquad=A_{\text{thursday}}\times3^{3}\]\[\qquad=345\times27\]\[\small\quad=\]